g(x) = ax + b / x, x≠k

given g(2)-3 & g(-2)=5


a) value of k

b) the values of a and b

c) the value of h if g-1(h)=2 (is this inverse)?


1
Expert's answer
2021-10-05T11:13:43-0400

Solution.


"g(x)=\frac{ax+b}{x}, x\neq k"

a) such as domain g(x) is equal "(-\infty,0)\bigcup (0,\infty)," so "k=0."

b) such as g(2)=-3, we will have "\frac{2a+b}{2}=-3." From here "2a+b=-6."

such as g(-2)=5, we will have "\frac{-2a+b}{-2}=5." From here "-2a+b=-10."

Solve system

"-2a+b=-10,\newline
2a+b=-6."

Solve this system by adding:"-2a+2a+b+b=-10-6,\newline
2b=-16,\newline
b=-16:2,\newline
b=-8."

Substitute the value of b in the first equation of the system, then

"-2a-8=-10,\newline
-2a=-2,\newline
a=1."

So, "a=1, b=-8."

c) "g(x)=\frac{x-8}{x}."

Find "g^{-1}(x)." To do this, express the variable x through the variable g:

"gx=x-8,\newline
gx-x=-8,\newline
(g-1)x=-8,\newline
x=\frac{-8}{g-1}=\frac{8}{1-g}."

And swap the variables g and x. We will have "g^{-1}(x)=\frac{8}{1-x}."

Such as "g^{-1}(h)=2," so

"\frac{8}{1-h}=2."

From here "1-h=4," and "h=-3."